The height of a male college freshman has a normal distribution with mean 71 inches and standard deviation 3 inches. What is the probability of selecting a student whose height is between 72 and 75 inches?
Answer #1

Solution :

Given that ,

mean = = 71

standard deviation = =3

P(72< x <75 ) = P[(72 - 71) / 3< (x - ) / < (75-71) /3 )]

= P( 0.33< Z < 1.33)

= P(Z <1.33 ) - P(Z <0.33 )

Using z table   

= 0.9082 -0.6293

probability= 0.2789